The radius of a circle is 5 cm and a chord of it is at a distance of 4 cm from the center. What is the length of the chord in cm?
Answer by jim_thompson5910(35256)

If you draw out the picture, or use the picture given by the book, then you'd see that a right triangle forms with legs of 'x' and 4. The hypotenuse is 5.

So
a = x
b = 4
c = 5

Let's use the pythagorean theorem to find x

a^2 + b^2 = c^2
x^2 + 4^2 = 5^2 ... plug in the given info; isolate x
x^2 + 16 = 25 ... square each value
x^2 + 16-16 = 25-16 ... subtract 16 from both sides
x^2 + 0 = 9
x^2 = 9
sqrt( x^2 ) = sqrt( 9 ) ... Apply the square root to both sides
x = 3

The leg of 3 represents exactly half of this chord.

So 3*2 = 6 is the length of the chord


Final Answer: 6 centimeters